A +4.5 μC charge is 22 cm to the right of a -8.4 μC charge. At the midpoint between the two charges what is the magnitude of the electric field?
Answer #1

The Electric field at midpoint = E = kq1/r^2 + kq2/r^2

E = (9*10^9/0.11^2)(4.5 + 8.4)*10^-6

E = 9.6*10^6 N/C